OpenStudy (anonymous):

HELP If you know the diameter of a circle, how do you find its circumference?

OpenStudy (anonymous):

\[circumference = \pi*diameter\]

OpenStudy (anonymous):

D'AWW THANKS BUD!(:

sam (.sam.):

\[circumference=2\pi(r)\] \[r=\frac{d}{2}\] \[circumference=2\pi(\frac{d}{2})\] \[circumference=\pi(d)\]